# Energy barriers for small electron polaron hopping in bismuth ferrite from first principles

### v1

@@ 3.2 Excess electrons at domain walls
When starting from a delocalized-electron configuration, a small electron polaron forms spontaneously in the plane labeled “1” (on the right-hand side of the wall, if the net polarization points to the right, see Figure 2(b)), which is a metastable site. The most stable site is inside the domain wall, in the plane labeled “0”.
@@
Figure 5 shows the energy barriers for electron hops at the 71° domain wall. Only the lowest barriers in each direction are shown. All barriers are depicted and tabulated in the Supplemental Information. The hops are projected on the symmetry-adapted coordinates (see Figure 5) [ wall, (a)], [within the wall, (b)], and [within the wall, , (c)]. For hops between the two trap states in the planes “0” and “1”, the barrier [bright symbols and lines in (a)] is similar to the thermal energy at room temperature, whereas for hops out of the wall (from plane “0” or “1” to plane “-1”,“2”, or “-2”) the barrier is about 2–3 times larger. For hops within the domain wall [(b) and (c)] there are pathes available with barriers similar to the thermal energy at room temperature. Barriers are invariant with respect to translation by a lattice vector parallel to the wall. Like in bulk, the barriers shrink upon compression. The lowest barriers for NN and 2NN hops are similar in size. The computed 2NN energy barriers are listed in Table 2 and in the Supporting Information.
double 2NN, | 2NN, | 2NN, escape | |
or | from DW | ||
% | 31 | 16 | 57 |
0 | 32 | 25 | 62 |
0.34% | 33 | 27 | 65 |
0.65% (extrapol.) | 35 | 29 | 67 |
The computed two-dimensional electron mobility inside the domain-wall plane is cm2/Vs, the computed two-dimensional diffusion constant inside the domain wall is cm2/s. and have an uncertainty of about one order of magnitude.
@@ 4 Discussion
Hall measurements for hematite resulted in electron mobilities of about 0.01 to 1 cm2/Vs [42, 43, 44], near the upper end of which lies the here computed mobility in bulk BiFeO3 of 0.6 cm2/Vs. Given the similarities in atomic and electronic structure of BiFeO3 and hematite, it seems plausible that intrinsic mobilities of BiFeO3 and hematite should be comparable.
@@
Experimentally determined activiation energies of electron conduction in BiFeO3 are typically much larger (between about 0.1 and 1.3 eV [45, 46, 47, 48, 49]) than the here computed barrier of about 28 meV, indicating that the experimentally determined barriers do not originate from electron polaron hopping in pristine bulk nor from escape of electron polarons from pristine neutral ferroelectric domain walls, assuming that the other typical walls, the 109° and the 180° walls, exhibit a similar or smaller trap depth than the 71° wall. It seems likely that the experimentally determined barriers of several hundred meV originate from defects other than pristine neutral ferroelectric domain walls, for example from point defects, such as oxygen or bismuth vacancies, or from interface effects, such as Schottky barriers. In Refs. [13, 50] an activation barrier of eV was found for -type conduction, which is about the energy difference between the electron polaron level and the conduction band minimum computed here. In principle, this activation barrier of 0.7 eV could originate from emission of electrons from the polaron state to the conduction band. This would mean that transport by emission from polarons into delocalized conduction states and subsequent ballistic transport would be more efficient than polaron hopping, inspite of the much higher energy barrier that must initially be overcome. However, there are reasons to believe that the lifetime of thermalized electrons in the conduction band at room temperature should be small. This is because the strong Fe- contribution to the states at the conduction band minimum and the small size of the electron polaron indicate that electron-phonon coupling should be large, which should lead to strong electron-phonon scattering. First-principles molecular dynamics simulations yield a polaron formation time of less than a picosecond [20]. Once a small electron polaron has formed, it is unlikely to transform into a delocalized conduction-band electron at room temperature since the energy distance between the polaron level and the conduction-band minimum of eV is much larger than the polaron hopping barrier of meV. It seems most likely that at least for thermalized electrons at temperatures room temperature, small electron hopping should be the dominant intrinsic -type transport process, and that in the reported experimental measurements of the electrical conductivity, the small barrier of meV for electron polaron hopping is hidden behind much larger barriers associated with emission from defects or interfaces other than pristine neutral ferroelectric domain walls.
@@ 5 Summary and conclusion
Energy barriers for small electron polaron hopping in BiFeO3 were determined using density-functional theory. Hopping in perfect, pristine BiFeO3 and at pristine, neutral ferroelectric 71° domain walls was considered. In both bulk and domain-wall plane, the calculated energy barriers are of the order of at room temperature ( meV) and are reduced by compressive strain. Using non-adiabatic Marcus theory, an electron mobility of cm2/Vs and cm2/Vs for bulk and within the domain-wall plane, respectively, is obtained at room temperature. The computed diffusion constant in bulk at room temperature is cm2/s both in the bulk and in the domain-wall planes. and have an uncertainty of about one order of magnitude. The domain wall planes act as traps for small electron polarons, with a trap depth of about two to three times at room temperature ( meV). These findings indicate that (a) the effect of pristine, neutral ferroelectric domain walls on the -type conductivity at room temperature consists in accumulating excess electrons and confining their motion to the domain-wall plane. In the Drude picture (conductivitycarrier densitymobility), the electron density along the wall is enhanced compared to the bulk, the mobility perpendicular to the wall is reduced compared to bulk. This is in line with experimental observations that conductivity is enhanced at ferroelectric domain walls and in directions parallel to the domain walls, (b) experimentally found activation barriers of several hundred meV for electrical conductivity in BiFeO3 should have an origin other than electron polaron hopping in pristine BiFeO3 or electron polaron escape from pristine neutral ferroelectric 71° domain walls.
@@ Acknowledgements
S. K. thanks Michele Reticcioli for recommending the OMC method. This project has received funding from a postdoc stipend of the University of Jena, Germany and from the Volkswagen Stiftung (Momentum) through the project “dandelion”. Computational resources and support were supplied by the HPC cluster ARA of the University of Jena, Germany. Figures were made using Vesta [51] and gnuplot.

### v2

@@ 3.3 Excess electrons at domain walls
3.3.1 Trapping of polarons at domain walls
The atomic structure, ionic ferroelectric polarization profile , total energy change with respect to a polaron position in the domain interior, relative polaron density as a function of temperature (: density in the domain interior), in-plane averaged and smoothened electronic potential , local direction-averaged strain, and octahedral tilt profiles around the three cartesian axes for small electron polarons near these domain walls are shown in Figure 6. The density was calculated with Eq. (1), the ionic polarization with Eq. (5).
The local strain with polaron in the lowest-energy site appears to couple to existing strain variations at the walls, possibly also to tilt variations. The local electronic potential in the pristine system, before polaron formation, exhibits a minimum at or near the lowest-energy polaron site.
The trapping energies for small electron polarons at the domain walls are listed in Table 4.
| DW | 160-atom supercell | 240-atom supercell |
| 71° | -0.063 | -0.051 |
| 109° | -0.075 | -0.065 |
| 180° | -0.044 | -0.054 |
We observe two different trap depths for the two 180° domain walls even though the two domain walls are equivalent by symmetry; this is because the symmetry operation relating first and second wall exchanges the magnetic sublattices. Here only polarons on spin-down sites were considered, the profile for polarons of the other spin was then obtained based on symmetry. The computed trapping energies (-51 meV for the 71°, -65 meV for the 109°, and -54 meV for the 180° domain wall) are similar to each other. The 180° domain wall differs from the 71° and 109° walls in that it attracts small polarons in a plane one layer apart from the layer closest to the wall, and, interestingly, each of the two equivalent 180° walls selectively traps essentially only polarons with a specific spin. The trapping energies were linearly corrected an for underestimated ferroelectric energy gain with PBEsol as the reference, see Table 1 and section 2.2. The short-range potential variations at the doman walls are superimposed on a large-scale potential landscape that depends on the boundary conditions, for example, open- or closed-circuit conditions, surface termination, etc. To estimate the density of excess electrons at the walls, one should consider both short-range and long-range potential. In a previous work [20], the long-range potential at 71° and 109° domain walls was found to consist in a potential step at the wall and a potential slope in the domains that depends on the boundary conditions. For zero potential slope (zero electric field in the domains), only the short-range potential acts on the polarons, and their density should be determined by Eq. 1, as depicted in Figure 6. For closed-circuit conditions, the potential step at the domain walls should create an electric field in the domains. This case is considered in Figure 7, assuming a domain-wall spacing of 150 nm [43, 44] and a polaron density in the dilute limit ( and very small).
3.3.2 Polaron hopping at domain walls
@@
For the 71° domain wall, polaron hopping was investigated along and across the domain wall. When starting from a delocalized-electron configuration at the 71° domain wall, a small electron polaron forms spontaneously in the plane labeled “1” (on the right-hand side of the wall, if the net polarization points to the right, see Figure 2(b)), which is a metastable site. The most stable site is inside the domain wall, in the plane labeled “0”.
@@
Figure 8 shows the energy barriers for electron hops at the 71° domain wall. Only some representative low barriers in each direction are shown. The computed 2NN energy barriers are listed in Table 5. All barriers are depicted and tabulated in the Supporting Information. The hops are projected on the symmetry-adapted coordinates [ wall, (a)], [within the wall, (b)], and [within the wall, , (c)]. For hops between the two trap states in the planes “0” and “1”, the barrier is about 0.2 eV; for hops out of the wall (from plane “0” or “1” to plane “-1”,“2”, or “-2”) the barrier is about 10 to 20% larger. The barriers are invariant with respect to translation by a lattice vector parallel to the wall. Like in the bulk, the barriers slightly increase under tensile strain. The lowest barriers for NN and 2NN hops are similar in size.
| double 2NN, | 2NN, | 2NN, escape | (cm2/Vs) | ||||
| or | from DW | (cm2/s) | |||||
| % | 0.19 | 0.19 | 0.22 | – | – | ||
| 0 | 0.20 | 0.19 | 0.23 | – | – | ||
| 0.34% | 0.20 | 0.20 | 0.23 | – | – | ||
| 1.9% (extrapol.) | 0.21 | 0.21 | 0.25 | ||||
The computed two-dimensional electron mobility inside the domain-wall plane is cm2/Vs, the computed two-dimensional diffusion constant inside the domain wall is cm2/s. Assuming 10% uncertainty in the barriers, one obtains 0.7 to 4 cm2/Vs and 0.2 to 1 cm2/s. The temperature dependence of the computed mobility is depicted in Figure 9. The mobility is dominated by the exponential decay originating in the hopping barrier.
@@ 4 Discussion
All three ferroelectric domain walls, the 71°, the 109°, and the 180° walls, trap small polarons on a short range, with a trap depth of about 50 to 65 meV. Depending on the boundary conditions, such as applied voltage or surface termination, a long-range potential exists in the case of the 71° and 109° domain wall, which originates from a potential step at the domain walls and can create an electric field in the domains, whose strength depends on the boundary conditions: the field should be maximal for closed-circuit conditions and zero for open-circuit conditions. For symmetric 180° domain wall, there should be no electric field. The electric field, if present, dominates the potential energy landscape for small polarons if the potential step at the wall is large compared to the trapping energy, which is the case for the 71° domain wall (potential step: 0.13 eV, trap depth: 51 meV). Assuming a resolution of 10 nm for spatially resolved conductivity measurements at domain walls, the increase in polaron density at the domain wall due to the short-range trap should be hidden behind the large density increase due to the long-range trap originating in the potential step at the wall. In the case of open-circuit conditions, the polaron density and conductivity increase at the wall should originate from the short-ranged trapping profile. In this case, only a moderate increase in polaron density and conductivity of about one order of magnitude is to be expected if measured on a ten-nanometer scale, and only at low temperatures, e.g. at 100 K. The electrical conductivity in the domain interior and at the domain walls will be further influenced by atomic point defects, which are not considered here.
Hall measurements for hematite resulted in electron mobilities of about 0.01 to 1 cm2/Vs [45, 46, 47], which is one to two orders of magnitude higher than the here computed mobility in bulk BiFeO3 of to cm2/Vs. Considering that the Fe-Fe distance in BiFeO3 is larger than in hematite, and that NN hops are spin-forbidden in BiFeO3, a lower mobility is to be expected in BiFeO3. In order to obtain the same mobility for BiFeO3 as in hematite, the energy barriers would need to be about 40% smaller, near 0.1 eV instead of 0.2 eV. At elevated temperatures, a combination of electron tunneling through the top of the barrier, smearing out of the traps due to thermal motion, incomplete self-trapping, and occupation of excited self-trapped states should effectively reduce the barriers. At low temperature, the effective barriers should approach those computed here. The here calculated hopping barriers ( eV) are similar to those obtained for PbTiO3 [31], so that the activation energy for the intrinsic electron mobility in BiFeO3 should be comparable to that in PbTiO3.
Experimentally determined activiation energies of electron conduction in BiFeO3 are typically in a similar range or larger (between about 0.1 and 1.3 eV [48, 49, 50, 51, 52]) than the here computed barrier of about 0.2 eV. The lower measured barriers could hence originate from electron polaron hopping, while it seems likely that the experimentally determined higher barriers of several hundred meV originate from defects other than pristine neutral bulk or ferroelectric domain walls, for example from point defects, such as oxygen or bismuth vacancies, or from interface effects, such as Schottky barriers.
@@
In Refs. [13, 53] an activation barrier of eV was found for -type conduction, which is about the energy difference between the electron polaron level and the conduction band minimum computed here. In principle, this activation barrier of 0.7 eV could originate from emission of electrons from the polaron state to the conduction band. This would mean that transport by emission from polarons into delocalized conduction states and subsequent ballistic transport would be more efficient than polaron hopping, inspite of the much higher energy barrier that must initially be overcome. However, there are reasons to believe that the lifetime of thermalized electrons in the conduction band at room temperature should be small: The strong Fe- contribution to the states at the conduction band minimum and the small size of the electron polaron indicate that electron-phonon coupling should be large, which should lead to strong electron-phonon scattering. First-principles molecular dynamics simulations yield a polaron formation time of less than a picosecond, see Supporting Information, similar to hematite [54]. For hematite, transient extreme-ultraviolet spectroscopy yielded “signatures characteristic of electron localization by small polarons […] within 100 fs following photoexcitation” [55]. Hence, it seems more likely that the measured barrier of 0.7 eV belongs to a defect or interface.
Once a small electron polaron has formed, it is unlikely to transform into a delocalized conduction-band electron at room temperature and below since the energy distance between the polaron level and the conduction-band minimum of eV is much larger than the polaron hopping barrier of eV. It seems most likely that small electron hopping should be the dominant intrinsic -type transport process at room temperature and below.
Here it is assumed that excess electrons have been brought into BiFeO3 in some way. For example, charge carriers may originate from point defects, such as oxygen vacancies or aliovalent dopants or impurities, from photoexcitation resulting in bound or dissociated electron-hole pairs, or from an electrical contact.
@@ 5 Summary and conclusion
Summarizing, the domain wall planes act as traps for small electron polarons; trapping energies of -51 meV (71° wall), -75 meV (109° wall), and -54 meV (180° wall) were obtained here. The origin of the trapping is likely a coupling of the polaron deformation to existing strain and possibly tilt variations at the domain walls. These trap depths are about two times at room temperature and hence near the boundary between deep and shallow levels. Assuming as the boundary between shallow and deep levels [56], the domain walls act as shallow traps at and above room temperature and become deep traps for temperatures below to K. The atomic and magnetic structure of the 180° domain walls couple in such a way that the walls selectively trap polarons depending on their spin. Pristine domain walls, especially 71° domain walls, can accumulate polarons in a region of tens of nanometers up to densities that are several orders of magnitude larger than in the bulk, depending on the boundary conditions. This is due to a potential step in the wall that can create a long-ranged electric field. For the 109° domain wall, a maximum polaron density of less than one order of magnitude more than in the bulk is possible based on the computed potential step, at symmetric 180° walls there should be no such long-range electric field. The experimentally measured increase in electrical conductivity at the domain walls should essentially originate in a local increase in polaron density at the walls.
For the first time, energy barriers for small electron polaron hopping in BiFeO3 were determined using density-functional theory. Also for the first time, electron polaron hopping at a ferroelectric domain wall was modeled from first principles. Hopping in perfect, pristine BiFeO3 and at pristine, neutral ferroelectric 71° domain walls was considered.
In both bulk and domain-wall plane, the calculated energy barriers for hopping are about 0.2 eV
and slightly increase under tensile strain.
The calculated hopping barrier for electrons in BiFeO3 is similar, and the resulting activation barrier for the electron mobility should be similar
to, e.g., PbTiO3. Using non-adiabatic Marcus theory, an electron mobility of
to cm2/Vs and a computed diffusion constant of 0.2 to 3 cm2/s
are obtained at room temperature both in the bulk and in the domain-wall planes.
Concluding, (a) experimentally found activation barriers of several hundred meV for electrical conductivity in BiFeO3
should have an origin other than electron polaron hopping in pristine BiFeO3 or electron polaron escape from pristine neutral ferroelectric 71° domain walls.
To identify the origin of these large energy barriers is here deferred to future work.
This will be a considerable task, but it should be undertaken to fully understand the limits of electronic conductivity in BiFeO3 and
other transition-metal oxides.
(b) The effect of pristine, neutral ferroelectric domain walls on the -type conductivity at room temperature consists in accumulating excess electrons
and confining their motion to the domain-wall plane.
In the Drude picture (conductivitycarrier densitymobility), the electron density at the wall is enhanced compared to the bulk,
the mobility perpendicular to the wall is slightly reduced compared to the bulk.
This is in line with experimental observations that conductivity is enhanced at ferroelectric domain walls
and in directions parallel to the domain walls.
At temperatures up to room temperature, pristine ferroelectric 71° domain walls can in principle
exhibit a carrier density and an electrical conductivity up to several orders of magnitude larger than in the bulk,
depending on the boundary conditions, due to a potential step of about 0.13 eV in the wall.
For the 109° domain wall, assuming a spatial resolution of ten nanometers,
a carrier density and conductivity enhancement by up to one order of magnitude is possible,
due to a potential step of about 17 meV.
For the symmetric 180° domain wall, the carrier density is enhanced by far less than one order of magnitude,
when averaged on a ten-nm scale.
Accumulation of point defects other than electron polarons, not considered here,
should further modify the potential and carrier density profile of the ferroelectric domain walls.
@@ Acknowledgements
S. K. thanks Michele Reticcioli for recommending the OMC method. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 746964, from a scholarship programme at the University of Jena within the Thuringian Scheme for Young Female Researchers and Young Female Artists funded by the Free State of Thuringia, and from the Volkswagen Stiftung (Momentum) through the project “dandelion”. Computational resources and support were supplied by the HPC cluster ARA of the University of Jena, Germany. Figures were made using Vesta [57] and gnuplot.